On January 1,2020 Versace Co. acquired 25% interest in the ordinary shares of Hayaan Mo Sila Inc. for ₱3,000,000 which reflected book value as of the date The Equipment has a remaining life of five years.


2020 2021

Net income ₱1,000,000 ₱1,500,000

Dividend declared 400,000 700,000

Assume the tax rate 30% and the depreciable asset is an ordinary asset (i.e., not subjected to final tax).

Questions:

Case No. 1: On January 3,2020 Versace Company sold an equipment costing ₱600,000 to Hayaan Mo Sila Company for ₱800,000.

1. Investment income 2020

   2. Investment 2021

Case No. 2: On January 3,2020 Hayaan Mo Sila Company sold an equipment costing ₱400,000 to Versace Company for ₱600,000.

3.Investment income in 2020

4. Investment income 2021



1
Expert's answer
2021-06-20T17:53:45-0400

Solution:

Case No. 1:

1.). Investment income in 2020:

Net income 2020 = 1,000,000

Share of net income = 1,000,000 "\times" 25"\%" = 250,000

Eliminate intra-entity profits from a downstream sale of equipment:

Gain from equipment sale = 800,000 – 600,000 = 200,000

Share of intra-entity profit = 200,000 "\times" 25"\%" = 50,000

Deduct this from the share of net income = 250,000 – 50,000 = 200,000

Investment income in 2020 = 200,000


2.). Investment income in 2021:

Net income 2021 = 1,500,000

Share of net income = 1,500,000 "\times" 25"\%" = 375,000

Investment income in 2021 = 375,000

 

Case No. 2:

3.). Investment income in 2020:

Net income 2020 = 1,000,000

Share of net income = 1,000,000 "\times" 25"\%" = 250,000

Eliminate intra-entity profits from the upstream sale of equipment:

Gain from equipment sale = 600,000 – 400,000 = 200,000

Share of intra-entity profit = 200,000 "\times" 25"\%" = 50,000

Deduct this from the share of net income = 250,000 – 50,000 = 200,000


Similarly add back depreciation expense = "\frac{50,000}{5} = 10,000"

Investment income = 200,000 + 10,000 = 210,000

Investment income in 2020 = 210,000

 

4.). Investment income in 2021:

Net income 2021 = 1,500,000

Share of net income = 1,500,000 "\times" 25"\%" = 375,000

Investment income in 2021 = 375,000